OpenStudy (anonymous):

anybody know how to do 2nd degree equations and systems

OpenStudy (anonymous):

Whats the question and i will tell you if i do or not.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

solve the equation:

OpenStudy (anonymous):

4x^2+9y^2=36 x^2-4y^2=16

OpenStudy (anonymous):

do u know to solve it so u could help me please

OpenStudy (anonymous):

Yeah i know how to do it. I will post the steps

OpenStudy (anonymous):

alright thank you

OpenStudy (anonymous):

is there certain way to solve this problem

OpenStudy (anonymous):

On second thought. Are you sure the second equation shouldnt read \[x^2 +4y^2 = 16\] Because i am getting imaginary numbers

OpenStudy (anonymous):

I was going to use the Elimination method but i am getting imaginary numbers when i do it that way.

OpenStudy (anonymous):

really

OpenStudy (anonymous):

the problem said

OpenStudy (anonymous):

solve the system: 4x^2+9y^2=36 x^2-4y^2=16

OpenStudy (anonymous):

what are the imaginary numbers you are getting?

OpenStudy (anonymous):

\[\sqrt{\frac{28}{25}}i\] I must be doing something wrong :/

OpenStudy (anonymous):

yea maybe because it said the answer has to be an ordered pair

OpenStudy (anonymous):

well you can write imaginary numbers in ordered pairs.. is this question in a math text book??

OpenStudy (anonymous):

or is it on a worksheet or something?

OpenStudy (anonymous):

its a computer program i have to do for homework

OpenStudy (anonymous):

how do you write imaginary numbers in an ordered pair?

OpenStudy (anonymous):

well thats what i was getting for the y value. So you would plug that into one of the original equations and solve for x and you would write it \[(x. \sqrt{\frac{28}{25}}i)\] But finding x would be a pain and i dont think that Y value is right to begin with.